OpenStudy (anonymous):

if have 2H2+C=CH4. 10 grams of hydrogen how much methane can you make

OpenStudy (4n1m0s1ty):

You first need to find the number of moles of hydrogen you have from 10 g.

OpenStudy (anonymous):

how do you do that 2H2+C=CH4

OpenStudy (4n1m0s1ty):

You need to know the molecular mass of hydrogen from the periodic table, then use it to convert to moles.

OpenStudy (anonymous):

I guess that would be 1.00794

OpenStudy (4n1m0s1ty):

Ok. I'm not sure how many sig figs your class wants, but thats about 1 g/mol for a hydrgen atom correct? So how many mols are in our sample?

OpenStudy (anonymous):

i have to use the balance equation which is 2H2+C=CH4. The actual question is if you have 10 grams of Hydrogen, how many methane can you make.

OpenStudy (4n1m0s1ty):

Once you have the number of mols in your sample you can answer that question.

OpenStudy (anonymous):

Im sorry, I am a home trying to help my daughter with her Chemistry. I have never had Chemistry and do not understand how to get the moles. I am totally loss here. .

OpenStudy (anonymous):

I a mom, sorry for typing home

OpenStudy (4n1m0s1ty):

Ok, I'll explain. Just give me a few seconds to write it up.

OpenStudy (anonymous):

thank you

OpenStudy (4n1m0s1ty):

The problem states that you have 10 g of hydrogen in your sample. We know that the molecular weight of hydrogen is 1 g/mol (I'm rounding to make the numbers easier). So to find the number of mols of hydrogen we just need to do some dimensional analysis.

OpenStudy (4n1m0s1ty):

We need to convert from grams to moles, so I'll setup a fraction like this: \[\frac{10\,g\,H}{}*\frac{1\,mol\,H}{1\,g\,H}= ?\,mols\,H\] This tells me that I need to divide the number of grams in by sample by the molecular weight in order to get the number of mols in my sample. By placing the grams unit on the bottom, it cancels out the grams unit on top of the other fraction, leaving the mole unit. This is probably not a good example since the hydrogen is only 1 g for each mol, so I'll show another example with carbon. \[\frac{60\,g\,C}{}*\frac{1\,mol\,C}{12\,g\,C}= 5\,mols\,C\] With carbon, you have a molecular weight of 12 g/mol (I"m using a sample of 60 g to make the numbers even). Here you can clearly see that the units cancel out,when I divide by the molecular weight.

OpenStudy (4n1m0s1ty):

I'll be back in a sec. Helping someone else out.

OpenStudy (anonymous):

ok

OpenStudy (4n1m0s1ty):

Ok, so was that explaination clear enough? About how to convert to number of mols from grams?

OpenStudy (anonymous):

yes, thank you

OpenStudy (anonymous):

I thought when looking at her textbook. You needed to use this formula, but was not sure.